You have been asked to forecast the additional funds needed (AFN) for Houston, Hargrove, & Worthington (HHW), which is planning its operation for the coming year. The firm is operating at full capacity. Data for use in the forecast are shown below. However, the CEO is concerned about the impact of a change in the payout ratio from the 10% that was used in the past to 30%, which the firm's investment bankers have recommended. Based on the AFN equation, by how much would the AFN for the coming year change if HHW increased the payout from 10% to the new and higher level? All dollars are in millions.

Last year's sales = S0 $300.0 Last year's accounts payable $50.0
Sales growth rate = g 40% Last year's notes payable $15.0
Last year's total assets = A0* $500 Last year's accruals $20.0
Last year's profit margin = PM 20.0% Initial payout ratio 10.0%

Select the correct answer.

a. $13.4
b. $11.7
c. $20.2
d. $15.1
e. $16.8
Answer #1

Calculate the AFN as follows:

AFN = (A0 / S0)*Change in sales - (L0 / S0)*Change in sales - M*S1* RR

Here,

A0 is the last year assets,

S0 is the last year sales,

L0 is the last year liability,

M is the profit margin,

S1 is the forecasted sales,

RR is the retention ratio.

------------------------------------------------------------------------------------

Forecasted sales = last year sales*(1+ growth rate) = 300*(1+40%) = 420

---------------------------------------------------------------

Change in sales = Forecasted sales - last year sales = 420 - 300 = 120

-----------------------------------------------------

When payout ratio is 30%.

Retention ratio =(1- Payout ratio) = (1-30%) = 70%

-----------------------------------------------------

AFN required when payout ratio is 30%

AFN = (500/300)*120 - ((50+20)/300)*120 - 20%*420*70%

AFN = 200 - 28 - 58.80

AFN = 113.20

---------------------------------------------

When payout ratio is 10%.

Retention ratio =(1- Payout ratio) = (1-10%) = 90%

-----------------------------------------------------

AFN required when payout ratio is 10%

AFN = (500/300)*120 - ((50+20)/300)*120 - 20%*420*90%

AFN = 200 - 28 - 75.60

AFN = 96.40

------------------------------------------

Calculate the increase in AFN as follows:

Increase in AFN = AFN when pay out ratio is 30% - AFN when payout ratio is 10%

Increase in AFN = $113.20 - $96.40

Increase in AFN = $16.80.